OpenStudy (anonymous):

What is the simplified form of 4z^2 - 4z - 15 over 2Z^2 + z - 15

sam (.sam.):

\[\frac{4z^2 - 4z - 15}{2z^2 + z - 15}\] factor \[\frac{(2z-5)(2z+3)}{(2z-5)(z+3)}\] \[\frac{(\cancel{2z-5)}(2z+3)}{\cancel{(2z-5)}(z+3)}\]

OpenStudy (anonymous):

Thank You (:

sam (.sam.):

welcome :)